OpenStudy (anonymous):

At a certain temperature, the solubility of strontium arsenate, Sr3(AsO4)2, is 0.0540 g/L. What is the Ksp of this salt at this temperature

OpenStudy (anonymous):

find the molar mass of the salt and dive .0540 by the molar mass and that is your kps

OpenStudy (anonymous):

i dont have a periodic table here, otherwise i would help you

OpenStudy (anonymous):

m.w.=540 so .0540/540 = 1*10-4=kps